OpenStudy (anonymous):

Write a problem based on the given information. b = price of a book 0.065b = tax 2(b + 0.065b) = 42.49 a.The sales tax at a bookstore is 6.5%. How many books did the customer buy if she spent $42.49? b.The sales tax at a bookstore is 6.5%. If a customer buys two of each book, how many different books can she buy with $42.49? c.The cost of two books is 42.49. What is the sales tax if the cost of one book is 0.065? d.The tax of a book is 0.065b whenever you buy two or more books. How many books can you buy for $42.49?

OpenStudy (anonymous):

I can help you

OpenStudy (anonymous):

ok thx, what do you think? i know it's not c

OpenStudy (anonymous):

Lets start with what do you think it is then? and why?

OpenStudy (kropot72):

@tori72 Do you agree that the price of a book plus the sales tax on the book is shown by (b + 0.065b) in the given equation?

OpenStudy (anonymous):

yea, i was thinking that it's choice B

OpenStudy (kropot72):

@tori72 Do you also agree that the amount of the tax, shown as '0.065b = tax' in the given information, represents the 6.5% sales tax on a book?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

you still there?

OpenStudy (kropot72):

Yes, I'm here. Can you now see the meaning of the given equation and briefly explain it? 2(b + 0.065b) = 42.49

OpenStudy (anonymous):

well if i distribute it i get 2b + 0.13 = 42.49 and solving for b gives me b = 21.18, so each book is 21.18

OpenStudy (anonymous):

wait hold on

OpenStudy (kropot72):

Correct. So you understand that the total cost of 2 books, including the sales tax on each of them, is $42.49.

OpenStudy (anonymous):

isn't it b = 9.93?

OpenStudy (kropot72):

Your equation 2b + 0.13 = 42.49 should be 2b + 0.13b = 42.49

OpenStudy (anonymous):

yea so 42.49/2.13 = 19.95

OpenStudy (kropot72):

Correct. But, once again, you appear to understand that the total cost of 2 books, including the sales tax on each of them, is $42.49. Is that correct?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

ok, i see it now, thanks!

OpenStudy (kropot72):

So what is your choice of the problem now?

OpenStudy (anonymous):

is it not b? a customer can buy 2 books if they have 42.49 and each book costs 21.18

OpenStudy (anonymous):

or a

OpenStudy (kropot72):

Option B says that the customer buys 2 of each book, meaning that if the customer buys only 2 books the books will be exactly the same title. Therefore to get 2 different books the customer would need to buy 4 books, 2 with the same title and 2 with another title.

OpenStudy (kropot72):

So your choice of option A is correct.

OpenStudy (anonymous):

thanks!

OpenStudy (kropot72):

You're welcome :)